OpenStudy (anonymous):

how do you simplify this without multiplying them all together? any help is appreciated (2x)^5 (5) (x^2+x+1)^4 (2x+1) + (x^2+x+1)^5 (5) (2x)^4 (2)

OpenStudy (anonymous):

convert into whole squares

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

sry try taking common like (2x)^4 out of the whole exp

OpenStudy (amistre64):

it looks maybe like binomial stuff

OpenStudy (dumbcow):

Factor out the greatest common factor: Both have a 5 in common Both have (2x)^4 in common Both have (x^2+x+1)^4 in common 5(2x)^4(x^2+x+1)^4[2x(2x+1) + 2(x^2 +x+1)]

OpenStudy (dumbcow):

then simplify, add like terms...

OpenStudy (amistre64):

im prolly wrong about that tho :)

OpenStudy (anonymous):

okay!! so then would you distribute the 2x and (2x+1) and then the 2 with the (x^2+x+1)??

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

so after you do that would you just take out the gcf out of that and thats it?

OpenStudy (dumbcow):

yep if you expand the (2x)^4 as well to 16x^4 you should get \[160x^{4}(x^{2}+x+1)^{4}(3x^{2}+2x+1)\]

OpenStudy (anonymous):

okay!! thank you so much! could i ask you one more..its a division problem and i want to make sure im doing it right

OpenStudy (dumbcow):

ok sure

OpenStudy (anonymous):

\[(2)(3-x)(2x)-x ^{2}(2)(-1)\div [2(3-x)]^{2}\] and obviously the denominator is supposed to be on the bottom

OpenStudy (anonymous):

hello?

OpenStudy (dumbcow):

well first you can simplify the top by multiplying 2 and 2x, 2 and -1 \[= \frac{4x(3-x) + 2x^{2}}{4(3-x)^{2}}\]

OpenStudy (dumbcow):

You can factor out a 2 on top resulting in \[\frac{2x(3-x) + x^{2}}{2(3-x)^{2}}\]

OpenStudy (anonymous):

can the (3-x) on top become a 1 and then the (3-x)^2 become just (3-x)?

OpenStudy (dumbcow):

the top can be then simplified to x(6-x) if you distribute and add like terms 6x-2x^2 +x^2 = 6x - x^2 = x(6-x) No they dont cancel because there is addition in the numerator

OpenStudy (dumbcow):

i think thats as far as you can go

OpenStudy (anonymous):

how did the 2x^2 just become x^2?

OpenStudy (dumbcow):

oh i added -2x^2 + x^2 = -x^2 -2 +1 = -1

OpenStudy (anonymous):

oh ok i didnt see that! and on this other expression \[(x+7)^{3}(4)(3x+1)^{3}(3) + 3(x+7)^{2}(3x+1)^{4}\] can you factor out that 3 or no because the 3 on the right is eventually distributed?

OpenStudy (dumbcow):

Yes it factors out it doesn't matter where the 3 is as long as its being multiplied by the rest of the term

OpenStudy (anonymous):

ok!!

OpenStudy (dumbcow):

\[3(x+7)^{2}(3x+1)^{3}(4(x+7)+(3x+1))\] \[=3(x+7)^{2}(3x+1)^{3}(7x+29)\]

OpenStudy (anonymous):

thats what i got! and for (x-5)-(x+7) / (x-5)^2 do you distribute the bottom?